Tohban Report 2009-03-04

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 25 Feb 2009
End Date: 04 Mar 2009
Tohban: Lindsay Glesener
Tohban email:
Next Tohban: Pascal Saint-Hilaire
List all reports


Solar Activity

Very quiet; no GOES flares and no new active regions.

Memory Management

Max SSR fill is 25-30%. The decimation is normal/active and front-segment events are being recorded during eclipse.

Data Gaps

On March 3, there is about half a day of sporadically missing data packets, likely from a bad pass. Since there was no solar activity during this time, it's probably not worth replaying the data.

Spacecraft Operations

On Feb. 26, the temperature threshold for the HV cutoff was changed from 100K to 115K, with yellow and red limits put in at 105K and 110K, respectively. This was done to ensure that we do not trip the cutoff since the temperature is now just below 100K. (CP2 is currently at 97.6K.) However, it does appear the the temperatures have recently been trending downward.

A HV change was requested: another 6 steps down on detectors 6 and 8.

Personal tools